OpenStudy (anonymous):

2.Suppose you roll an eight-sided die two times hoping to get two numbers whose sum is even. What is the sample space? How many favorable outcomes are there?

OpenStudy (anonymous):

i dont understand this part - two dimensions, each ranging from 1 to 8

OpenStudy (anonymous):

A sample space is a set consisting of all the possible outcomes of an event. We use brackets curly brackets to show the items in a set.

OpenStudy (kirbykirby):

i'm ok with your sample space, but maybe also specify that \(x,y\in \mathbb{Z}\)

OpenStudy (anonymous):

I don't know my brother helped me.

OpenStudy (anonymous):

guess he's wrong

OpenStudy (kirbykirby):

32 sounds right for the favourable outcomes

OpenStudy (anonymous):

when rolling an eight sided die it ranges {1 2 3 4 5 6 7 8 }

OpenStudy (kirbykirby):

{\((x,y)|1\le x,y\le8, x,y\in \mathbb{Z}\)}

OpenStudy (anonymous):

why don't we just ignore that and start over. I'll give you some formulas.

OpenStudy (kirbykirby):

its the same sample space you provided but with an extra condition... this is to make sure you use the values 1,2,3,...,8 only and not a continuous range (1,8)

OpenStudy (anonymous):

Sample space = 8x8 = 64 Favorable outcomes = 32

OpenStudy (anonymous):

kitby i think that in ur sample space there are sums of number whose result is even and odd too

OpenStudy (kirbykirby):

the same space is all possibilities you can encounter. The favourable outcomes are only ones for which the two numbers added together are even

OpenStudy (anonymous):

oh! i got it it is the sample space 64

OpenStudy (anonymous):

i was thinking in the favorable outcomes

OpenStudy (kirbykirby):

Yeah there are 64 possibilities. But we usually enumerate all poissiblties when they ask for "the sample space", which is what @Martinmelissa was trying to do writing it in set notation

OpenStudy (anonymous):

if there is 32 favorable outcomes how can you put it in a sample space? he is asking for two numbers whose sum is even.

OpenStudy (kirbykirby):

We want the number of favourable outcomes. So this ok the pairs you get for which their sum is even: (1,1), (1,3), (1,5), (1,7) (2,2), (2, 4), (2, 6), (2, 8) ....

OpenStudy (anonymous):

ok so the sample space would be what?

OpenStudy (anonymous):

since the two numbers are even.

OpenStudy (kirbykirby):

yeah, with all the pairs enumerated

OpenStudy (kirbykirby):

well the sum of the 2 numbers is even

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I really need help. I don't get this. Sample space is hard.

OpenStudy (whpalmer4):

sum of two odd numbers is also even, don't forget

OpenStudy (whpalmer4):

it's 1 even + 1 odd that gives you an odd number. it may be helpful to count the odd outcomes and subtract them from the total outcomes to get the even outcomes, rather than trying to determine directly...

OpenStudy (kirbykirby):

The sample space is the list of \(all\) possible outcomes when you roll the two dice, regardless of what the sum of the 2 numbers is. So, we get that: S={\((1,1),(1,2),(1,3),...,(1,8),\) \((2,1),(2,1),(2,3),...,(2,8),\) \(...\) \((8,1),(8,2),(8,3),...,(8,8)\)}

OpenStudy (anonymous):

32- the odd outcomes = ....

OpenStudy (kirbykirby):

this will give 8 rows, with 8 pairs (x,y) in each row... which gives 8*8=64 outcomes in the sample space Now you can list the favourable outcomes in a systematic way. Use 1 row to count all possibilities for the 1st die: (1,1), (1,3), (1,5), (1,7) (2,2), (2, 4), (2, 6), (2, 8) (3,1), (3, 3), (3, 5), (3, 7) ... (8,2) ,(8,4), (8,6), (8,8) You can see that when he 1st die is odd, there are 4 possibilities . So, when the first die is 1,3,5, or 7, we get 4 possibilities for each. So 4*4=16 We get the same thing when the 1st die is even: when the 1st die is 2,4,6,8 we still get 4 valid pairs for each number. So 4*4 = 16 So in total there are 16+16 = 32 favourable outcomes

OpenStudy (kirbykirby):

You don't have to actually list all the favourable pairs. You should notice a clear pattern that allows you to deduce the information above^^

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

is this right? it is another question. just making sure Pablo rolled a standard die sixty times. He got a 1 twelve times. How does this result compare to the expected results? Explain. Since a die has 6 sides, one expects each number to be rolled 1/6 of the time. For 60 rolls, one would expect 60(1/6) = 10 rolls to in 1.12 rolls of 1 is 2 more than expected.

OpenStudy (kirbykirby):

i'm not sure I get the 10 rolls in 1.12 rolls part. But Other than that it seems good. He would expect a "1" 10 times, but he got it in 12 times, which is 2 more rolls than what's expected

OpenStudy (anonymous):

thanks

OpenStudy (kirbykirby):

your welcome